What distance separates two objects of equal mass (5 kg) that are experiencing a gravitational force of 100 N?



1
Expert's answer
2022-04-26T14:54:59-0400

"F=\frac{Gm^2}{r^2},\implies"

"r=m\sqrt\frac GF=20.4~\mu m ."